Directions: Find the equation of the line that passes through each of the following pairs of points.


1) (1,7) and (4,22) 2) (-2,13) and (2,3) 3) (0,-10) and (16,2)

PS: These next problems are fractions and I am going to use a "/" to seperate them. The first one is 3 and negative one and a half. The last set is one comma one and three secondths.

4) (3, -1/2) and (1, 1 3/2)

Thank you! I have math soon!
1
Expert's answer
2019-11-11T09:37:42-0500

A(x1;y1) B(x2;y2)

"(x-x1)\/(x2-x1)=(y-y1)\/(y2-y1)"

1)A(1,7) B(4,22);

(x-1)/(4-1)=(y-7)/(22-7);

Answer: (x-1)/3=(y-7)/15.

2)A(-2,13) B(2,3);

(x+2)/(2+2)=(y-13)/(3-13).

Answer: (x+2)/4 = (y-13)/(-10).

3)A(0,-10) B(16,2);

(x-0)/(16-0)=(y+10)/(2+10).

Answer: x/16=(y+10)/12.

4)A(3,-1/2) B(1, 1 3/2);

(x-3)/(1-3) = (y+1/2)/(2.5+0.5).

Answer: (x-3)/(-2)= (y+0.5)/3.